# A Dataset for Large Language Model-Driven AI Accelerator Generation

### v1

@@ 6 Conclusion
This study has introduced the first publicly accessible LLM prompt-Chisel code dataset, dubbed SA-DS. The prompt and code examples in SA-DS cover wide variety of applications and design criteria. A proof-of-concept experiment has showcased the benefits of SA-DS in enabling the high-quality generation of hardware accelerator designs with mere verbal descriptions of novice users. This exemplifies the promising potential of enabling further research in the area of utilizing LLMs for automated hardware design generation. Key examples along this line include fine-tuning high-end LLMs for hardware design, optimized multi-shot learning, and prompt engineering serving the objectives of design efficiency in terms of execution time, hardware cost, and power consumption.

### v2

@@ V Conclusion
This study has introduced the first publicly accessible LLM prompt-Chisel code dataset, dubbed SA-DS. The prompt and code examples in SA-DS cover a wide variety of applications and design criteria. A proof-of-concept experiment has showcased the benefits of SA-DS in enabling the high-quality generation of hardware accelerator designs with mere verbal descriptions of novice users. This exemplifies the promising potential of enabling further research in the area of utilizing LLMs for automated hardware design generation. Key examples along this line include fine-tuning high-end LLMs for hardware design, optimized multi-shot learning, and prompt engineering serving the objectives of design efficiency in terms of execution time, hardware cost, and power consumption.